XEM: Notable Events and Changes

From HallCWiki
Jump to: navigation, search

October 19, 2022

  • BCM 3H05 (Unser girder) temperature out of control (https://logbooks.jlab.org/entry/4063487)
    • Temperature controller was replaced - no data taken with bad temperature I think
    • Later, observed occasional temperature excursions up to 113. Not sure if this is a problem.
      • Shuji discovered this is completely correlated with target motion!

October 14-15, 2022

October 15, 2022

  • SHMS trigger changes
    • ELLO = 3/4 AND Cherenkov (was 3/4 AND Cherenkov AND PRLO)
    • ELHI = 3/4 AND PRLO (was 3/4 AND PRHI)

October 10, 2022

  • SHMS 2Y- scalers restored during the middle of SHMS run 17336 - those scalers were missing for all runs before this
  • Noticed spurious signal in SHMS 2Y21- - it turned out to be RF signal plugged into FADC, and map fiule wasn't updated. Commented this out in the map file - does not impact data. Could impact (maybe?) data analyzed before map file change

October 9, 2022

October 8, 2022

  • SHMS NGC PMT1 HV changed back to 1700 V

October 7, 2022

October 6, 2022

  • SHMS PRLO and PRHI thresholds increased to -45.4 and -50.9 mV to get rid of huge "no beam rate."
    • These thresholds were used for most of SHMS paddle study - affects runs starting with 17261
  • Mark updated the trigger mapping for SHMS DAQ (see elog https://logbooks.jlab.org/entry/4051681)
    • Would only have impacted data taken with elclean trigger - we had not taken any to this point, so not an issue

October 4 (day)-9 (swing), 2022

  • No beam due to magnets connected to RSEP8 power supply - magnets were overheating due to inadequate LCW flow

October 4, 2022

October 3, 2022